# The Need for Benchmarks to Advance AI-Enabled Player Risk Detection in Gambling

### v1

@@ 3.2 Types of Benchmarking
Benchmark datasets are as varied as ML itself and can be categorized along several axes. At a fundamental level, they can be distinguished by data type: images, audio, text, time-series, and tabular data. They can also be characterized by the ML task they address: classification, detection, regression, etc. The data curation method can also characterize benchmarks, with datasets ranging from purely synthetic data (e.g., see ACT-Thor: hanna-etal-2022-act) to real-world data (e.g., SVHN).
Beyond these characteristics, benchmarking can be categorized as scientific-based or competition-based. As described by dueben2022challenges, scientific benchmarks provide a platform for addressing broad, long-term problems shared by many research groups, with performance often self-reported through peer-reviewed publications. In contrast, competition benchmarks generally focus on narrower tasks that benefit from short-term, intensive focus of a wider community. Kaggle is a well-known example where monetary awards are often given for the top-performing solution. For instance, the “American Express – Default Prediction” competition challenged participants to use anonymized transactional data to predict which customers would fail to pay their debts (amex-default-prediction). Other platforms, such as HuggingFace, Papers with Code, the UCI ML Repository, and OpenML, also facilitate benchmarking (typically without monetary components) by providing infrastructure to host data, share models, and track performance on public leaderboards. As such, the line between these categories can blur.
@@ 3.3 The State of Benchmarking in Player Risk Detection
Despite the success of benchmarking in other fields, its application in gambling studies remains nascent, with no formal benchmark datasets for player risk detection. The closest example is research stemming from The Transparency Project [^7] For details on this initiative, see The Transparency Project’s website: http://www.thetransparencyproject.org/, which made several datasets publicly available. The bwin dataset, in particular, functioned as a de facto benchmark for researchers interested in player risk detection. The bwin dataset consists of player account records from a European-based online sports betting provider. A subset of these accounts is labeled with the reason for account closure, including “gambling problems,” which commonly serves as a proxy for problematic play. This dataset enabled foundational ML studies, from initial cluster analyses (adami2013markers; braverman2012gamblers) to later classification tasks comparing various supervised ML models (Philander2013). However, the dataset’s utility is limited by its age and the fact that it represents data from only a single operator. Although ML studies have continued to emerge since The Transparency Project, these efforts remain fragmented. Research groups often produce valuable work, but it frequently relies on proprietary data that cannot be publicly shared for broader validation (Ghaharian2022).
@@
Thus, a primary barrier to establishing benchmarks in gambling studies is data accessibility. Because player data is commercially sensitive and proprietary, it creates perceived challenges in data sharing and research transparency. Although early initiatives like The Transparency Project partially overcame these issues, their impact was limited by the technological constraints of the time. More recently, the University of Nevada, Las Vegas (UNLV) Payments Research Collaborative adopted a more modern data-sharing model, publishing a peer-reviewed paper on a dataset of gambling payment transactions and making analytic subsets available on a public repository (ghaharian2023payments). Similarly, Zendle2024 contributed by releasing a synthetic dataset [^8] A “synthetic” dataset is an artificially generated dataset. In this case, the authors used their original dataset to generate a new dataset that mimics the statistical patterns and correlations of the original. This process allows researchers to share data for analysis without exposing any of the original, privacy-sensitive data. derived from the banking transaction data of 424 UK-based survey respondents, linked with their Problem Gambling Severity Index (PGSI) scores. Still, without operator cooperation, these efforts are limited in their generalizability.
@@ 4.1 Core Dimensions
Time refers to the length of the period of play that the data spans. For explanatory purposes in this paper, we sub-categorize time into sessions, days, and months. Providing a short-term view (i.e., sessions) is crucial for identifying “in-the-moment” risk behaviors, such as within-session loss-chasing (Parke2015). Longer-term horizons, such as days or months, are necessary to understand cumulative impacts and identify at-risk players within specific time-frames. For example, prior studies have sought to assess future risk using as little as 7 days (auer2024predicting) or 30 days of data (braverman2012gamblers), while others have used 12-month datasets to create ML models (kairouz2023enabling; Murch2023).
@@
Gambling Vertical refers to the type of gambling products offered. The main gambling verticals by market share are lottery, casino, and sports betting. Within these verticals there are sub-categories—for example, table games, slots, and live dealer in casino, draw products and instant win games in lottery, and fixed odds, prop, and parlay in sports betting. While ML models can be developed to encapsulate player activity across gambling verticals, prior work has suggested superior performance when vertical-specific models are developed (e.g., see kairouz2023enabling). Thus, a benchmarking suite should strive to provide datasets that can be used to develop both broad and vertical-specific models.
@@ 4.2 Example Benchmark Datasets
Importantly, a single benchmark dataset can be used to support multiple benchmark tasks. A task is defined as the specific ML exercise a participant must complete. For instance, a dataset containing PGSI scores as the target variable could support multiple tasks. These might include a binary classification problem, where the goal is predicting whether a player is at risk based on PGSI scores (e.g., 0 = no risk, 5+ = at-risk), or a multi-class classification problem, where the goal is predicting risk across multiple PGSI buckets (e.g., 0, 1-2, 3-4, 5-7, 8+).
@@ 4.3 Requirements
As dueben2022challenges describe, “Benchmark datasets are defined by the community. While many datasets can be proclaimed as ‘benchmarks,’ they are only successful if the ML community embraces them.” Similarly, the success of a benchmarking suite for player risk detection hinges on its adoption by relevant stakeholders. As a foundation for future implementation of such a benchmarking suite, we present initial discussion on data specifications as well as various technical and usability requirements. Many of the administrative requirements could be satisfied by existing data repositories (for in-depth discussion see: longjohn2024benchmark) and could be worked into a customized repository specific for this task domain.
@@ 4.3.2 Dataset documentation
Because raw data are rarely ready for immediate use, they require processing to become a usable benchmark dataset. The benchmarking suite must facilitate recording of any processing steps taken, and given the sensitivity of player protection – where model performance directly impacts human well-being – there should also be an attempt to enforce quality review. As highlighted by longjohn2024benchmark, there is currently a lack of quality review of datasets, and existing repositories often view themselves as simply providing infrastructure to support data sharing rather than providing review or quality assurance processes. To address this, the proposed benchmarking suite could require the publication of peer-reviewed introductory papers to accompany datasets, similar to those found in Data in Brief (e.g., see: ghaharian2023payments).
The entire suite would house numerous datasets. Thus, users must understand each dataset’s context and structure. Best practices from established data science platforms and repositories can provide a strong foundation (e.g., HuggingFace, UCI ML Repository, OpenML). For instance, datasets should include core metadata such as the size and shape of the data, list a creator/owner of the dataset (who can act as a point of contact for queries and/or issues), as well as versioning and timestamps. Some datasets may well reside in raw format (e.g., session-level activity for each bet placed), whereas others may represent aggregated information (e.g., session-level activity rolled up to daily or weekly information about players’ activity). Tasks associated with benchmark datasets could also be listed alongside this metadata and/or with further details about each task (e.g., train-test splits, primary evaluation metric, etc.) provided in a separate location within the repository (e.g., as a downloadable file or a link to a dedicated page). We provide an example of such metadata for the Early Detection Benchmark Dataset (example D above) in Table 1.
@@
This “data work” is a substantial undertaking but is essential to the downstream success of any benchmark dataset. Unfortunately, this type of work is often undervalued and under-incentivized (gero2023incentive). To incentivize and recognize these efforts, datasets could be treated as citable intellectual contributions through persistent identifiers, such as a Digital Object Identifier (DOI) (longjohn2024benchmark).
@@ 5.1 Governance and Administration
A third-party model involving a neutral entity like an international research institute or data trust presents a distinct advantage in this regard (Paprica2020). Such an organization is not constrained by jurisdictional borders and could therefore aggregate data across multiple states, provinces, and nations. This transnational approach would enable a far more comprehensive dataset.
@@ 5.2 Dataset Size
The use of synthetic data provides a potentially constructive path to scale dataset size while avoiding issues with player privacy and commercial data sharing sensitivity (samsel2025). By generating an artificial dataset that replicates the statistical properties of real-world data without containing any personally identifiable information, synthetic data allows for broad access for research and model testing while the source data remains secure. The viability of this approach, however, is entirely contingent upon the quality and representativeness of the original dataset from which the synthetic version is derived, as well as the assumptions of the approach used to generate the synthetic data (Chen2021; Susser2024).
@@ 5.3 Defining the Target Variable
First, proxy indicators of harm such as account closure and, more commonly, voluntary self-exclusion (VSE) are used by both industry and academic studies (e.g., finkenwirth2021using; hopfgartner2023predicting; Percy2016). These are routinely collected alongside the data used to make predictions, can easily delineate the at-risk group through binary classification (VSE or no VSE), and are not subject to self-report bias or measurement error. However, their reliability as indicators of at-risk status is questionable. Players can choose to self-exclude or close their accounts for various reasons beyond the experience of gambling problems, including to pre-empt harms or to close one’s online account (catania2021understanding; auer2016should). One strategy to mitigate this risk is to use a subset of self-excluders who have a sufficiently long track record of play with the operator, to ensure players who self-exclude relatively quickly after opening an account are not used for model training.
@@
Second, screening instruments such as the PGSI (ferris2001canadian) or Brief Biosocial Gambling Screen (gebauer2010optimizing) have been used as target variables in recent academic studies (auer2023using; hopfgartner2024using; kairouz2023enabling; Murch2023; Perrot2022) and by Focal Research Consultants (hancock2008gambling; Schellinck2011). Self-report measures offer a more reliable indicator of risk status, although are more difficult to obtain and link with the data used to make predictions, are subject to self-report biases (particularly if administered by the operator), and require a threshold for classification as at-risk to be selected in advance (Murch2023).
@@
Third, a final approach may be the use of single indicators such as loss chasing or substantial variations from normal play, which are themselves sometimes used to classify consumers as at-risk in the absence of ML (McAuliffe2022). Several researchers have attempted to develop thresholds for standard behavioral indicators (e.g., frequency of gambling, amount spent) that differentiate lower- from higher-risk gambling to form lower risk gambling limits (hodgins2023lower; louderback2021developing) – these, or similar threshold-style indicators, could plausibly serve as targets, either individually or in combination.
@@
Overall, the variation and lack of a standard approach to classification presents a challenge to constructing and comparing ML models, as there is no agreed-upon specification to label training data appropriately. Given the diversity of outcome variables and definitions of risk, we do not propose a single ground truth. Instead, we recommend a suite of benchmarks that can evolve over time as the domain continues to develop and as a clearer definition of “at-risk” emerges. We do, however, advocate for transparency in how at-risk status is determined to ensure the process can be replicated in real-world environments. For example, human-based labeling of cases based on opaque reasoning would not satisfy this criterion, but labels determined using a clear and reproducible set of steps would.
@@ 5.4 Demographic and Jurisdictional Diversity
Another challenge is significant diversity in player populations and jurisdictions. A successful benchmarking framework must avoid over-representing datasets from one demographic group (e.g., all men) or market (e.g., North America), except where the benchmark results are explicitly only applied within those settings. Over-representing certain demographic groups in benchmark datasets risks validating the performance of models that perform poorly in real-world settings with more diverse populations and could potentially favor biased models that disproportionately target or ignore minority groups (see Murch2024).
There is evidence that ML models can effectively detect at-risk gambling in datasets from countries not included in training data, although training on local data can improve performance (hopfgartner2024using). A fairer benchmarking framework will offer datasets from a variety of regions to allow contributors to demonstrate the performance of their model(s) in preferred settings (e.g., EU vs. North America).
@@ 5.5 Dataset Characteristics
Jurisdictional variation may also impact the ease with which models are applied due to differences in feature availability. For example, a model built using online tracking data in Australia where players can cancel their withdrawals may attribute high importance to this behavior, which would not be present in datasets from jurisdictions that have prohibited this feature (e.g., Great Britain; gamblingcommission2021consultation). This highlights the need for clear documentation that describes the characteristics of each benchmarking dataset and idiosyncrasies resulting from regulation or product innovation.
@@
A further consideration is the temporal structure of the data. Academic studies report building models using varying timeframes of tracking data, ranging from 12 months (e.g., Murch2023; Perrot2022) to just 30 days (hopfgartner2024using). As discussed in Section 4, there is a need to provide benchmarks for different periods that span months to days. Ideally, benchmark datasets will span an extended period (e.g., 12 months) to permit varied testing on the same dataset and the use of diverse timeframe aggregations (e.g., monthly or daily behavioral averages). A benchmark for acute risk detection (e.g., 1 to 7 days) may benefit from uniquely large datasets, since these can include all active and newly registered customers, not only those with extensive historical data.

### v2

@@ 3.2 Types of Benchmarking
Benchmark datasets are as varied as ML itself and can be categorized along several axes. At a fundamental level, they can be distinguished by data type: images, audio, text, time-series, and tabular data. They can also be characterized by the ML task they address: classification, detection, regression, etc. The data curation method can also characterize benchmarks, with datasets ranging from purely synthetic data (e.g., see ACT-Thor: Hanna et al., 2022) to real-world data (e.g., SVHN).
Beyond these characteristics, benchmarking can be categorized as scientific-based or competition-based. As described by Dueben et al. (2022), scientific benchmarks provide a platform for addressing broad, long-term problems shared by many research groups, with performance often self-reported through peer-reviewed publications. In contrast, competition benchmarks generally focus on narrower tasks that benefit from short-term, intensive focus of a wider community. Kaggle is a well-known example where monetary awards are often given for the top-performing solution. For instance, the “American Express – Default Prediction” competition challenged participants to use anonymized transactional data to predict which customers would fail to pay their debts (Howard et al., 2022). Other platforms, such as HuggingFace, Papers with Code, the UCI ML Repository, and OpenML, also facilitate benchmarking (typically without monetary components) by providing infrastructure to host data, share models, and track performance on public leaderboards. As such, the line between these categories can blur.
@@ 3.3 The State of Benchmarking in Player Risk Detection
Despite the success of benchmarking in other fields, its application in gambling studies remains nascent, with no formal benchmark datasets for player risk detection. The closest example is research stemming from The Transparency Project [^8] For details on this initiative, see The Transparency Project’s website: http://www.thetransparencyproject.org/, which made several datasets publicly available. The bwin dataset, in particular, functioned as a de facto benchmark for researchers interested in player risk detection. The bwin dataset consists of player account records from a European-based online sports betting provider. A subset of these accounts is labeled with the reason for account closure, including “gambling problems,” which commonly serves as a proxy for problematic play. This dataset enabled foundational ML studies, from initial cluster analyses (Adami et al., 2013; Braverman and Shaffer, 2012) to later classification tasks comparing various supervised ML models (Philander, 2013). However, the dataset’s utility is limited by its age and the fact that it represents data from only a single operator. Although ML studies have continued to emerge since The Transparency Project, these efforts remain fragmented. Research groups often produce valuable work, but it frequently relies on proprietary data that cannot be publicly shared for broader validation (Ghaharian et al., 2022).
@@
Thus, a primary barrier to establishing benchmarks in gambling studies is data accessibility. Because player data is commercially sensitive and proprietary, it creates perceived challenges in data sharing and research transparency. Although early initiatives like The Transparency Project partially overcame these issues, their impact was limited by the technological constraints of the time. More recently, the University of Nevada, Las Vegas (UNLV) Payments Research Collaborative adopted a more modern data-sharing model, publishing a peer-reviewed paper on a dataset of gambling payment transactions and making analytic subsets available on a public repository (Ghaharian et al., 2023). Similarly, Zendle and Newall (2024) contributed by releasing a synthetic dataset [^9] A “synthetic” dataset is an artificially generated dataset. In this case, the authors used their original dataset to generate a new dataset that mimics the statistical patterns and correlations of the original. This process allows researchers to share data for analysis without exposing any of the original, privacy-sensitive data. derived from the banking transaction data of 424 UK-based survey respondents, linked with their Problem Gambling Severity Index (PGSI) scores. Still, without operator cooperation, these efforts are limited in their generalizability.
Regulator-coordinated requests have proven effective for targeted research: in the UK, the Gambling Commission worked with GambleAware to commission player account data across seven major operators, although barriers limited the availability of the data to third party researchers (Forrest and McHale, 2022). Additionally, legislative mandates could offer a more structural approach. For example, the Massachusetts Expanded Gaming Act requires casino licensees to provide anonymized customer tracking data to the Massachusetts Gaming Commission, which must make this data available to qualified researchers (Andrews et al., 2025).
@@ 4.1 Core Dimensions
Time refers to the length of the period of play that the data spans. For explanatory purposes in this paper, we sub-categorize time into sessions, days, and months. Providing a short-term view (i.e., sessions) is crucial for identifying “in-the-moment” risk behaviors, such as within-session loss-chasing (Parke et al., 2015). Longer-term horizons, such as days or months, are necessary to understand cumulative impacts and identify at-risk players within specific time-frames. For example, prior studies have sought to assess future risk using as little as 7 days (Auer and Griffiths, 2024) or 30 days of data (Braverman and Shaffer, 2012), while others have used 12-month datasets to create ML models (Kairouz et al., 2023; Murch et al., 2023).
@@
Gambling Vertical refers to the type of gambling products offered. The main gambling verticals by market share are lottery, casino, and sports betting. Within these verticals there are sub-categories—for example, table games, slots, and live dealer in casino, draw products and instant win games in lottery, and fixed odds, prop, and parlay in sports betting. While ML models can be developed to encapsulate player activity across gambling verticals, prior work has suggested superior performance when vertical-specific models are developed (e.g., see Kairouz et al., 2023). Thus, a benchmarking suite should strive to provide datasets that can be used to develop both broad and vertical-specific models.
@@ 4.2 Example Benchmark Datasets
Importantly, a single benchmark dataset can be used to support multiple benchmark tasks. A task is defined as the specific ML exercise a participant must complete. For instance, a dataset containing PGSI scores as the target variable could support multiple tasks. These might include a binary classification problem, where the goal is predicting whether a player is at risk based on PGSI scores (e.g., 0 = no risk, 5+ = at-risk), or a multi-class classification problem, where the goal is predicting risk across multiple PGSI buckets (e.g., 0, 1-2, 3-4, 5-7, 8+). However, acknowledging that validated self-report measures are not always available, tasks could also be defined around other harm proxies (See Section 5.3 for a discussion).
@@ 4.3 Requirements
As Dueben et al. (2022) describe, “Benchmark datasets are defined by the community. While many datasets can be proclaimed as ‘benchmarks,’ they are only successful if the ML community embraces them.” Similarly, the success of a benchmarking suite for player risk detection hinges on its adoption by relevant stakeholders. As a foundation for future implementation of such a benchmarking suite, we present initial discussion on data specifications as well as various technical and usability requirements. Many of the administrative requirements could be satisfied by existing data repositories (for in-depth discussion see: Longjohn et al., 2024) and could be worked into a customized repository specific for this task domain.
@@ 4.3.2 Dataset documentation
Because raw data are rarely ready for immediate use, they require processing to become a usable benchmark dataset. The benchmarking suite must facilitate recording of any processing steps taken, and given the sensitivity of player protection – where model performance directly impacts human well-being – there should also be an attempt to enforce quality review. As highlighted by Longjohn et al. (2024), there is currently a lack of quality review of datasets, and existing repositories often view themselves as simply providing infrastructure to support data sharing rather than providing review or quality assurance processes. To address this, the proposed benchmarking suite could require the publication of peer-reviewed introductory papers to accompany datasets, similar to those found in Data in Brief (e.g., see: Ghaharian et al., 2023).
The entire suite would house numerous datasets. Thus, users must understand each dataset’s context and structure. Best practices from established data science platforms and repositories can provide a strong foundation (e.g., HuggingFace, UCI ML Repository, OpenML). For instance, datasets should include core metadata such as the size and shape of the data, list a creator/owner of the dataset (who can act as a point of contact for queries and/or issues), as well as versioning and timestamps. Some datasets may well reside in raw format (e.g., session-level activity for each bet placed), whereas others may represent aggregated information (e.g., session-level activity rolled up to daily or weekly information about players’ activity). Critically, documentation should specify the temporal relationship between the player activity data and any associated labels (the target variable). This includes defining the timeframe the activity data covers (e.g., the first 7 days of play, 12-months since registration date) and its temporal proximity to the target variable (e.g., a PGSI score collected at 6 months, the date self-exclusion occurred). Tasks associated with benchmark datasets could also be listed alongside this metadata and/or with further details about each task (e.g., train-test splits, primary evaluation metric, etc.) provided in a separate location within the repository (e.g., as a downloadable file or a link to a dedicated page). We provide an example of such metadata for the Early Detection Benchmark Dataset (example D above) in Table 1.
@@
This “data work” is a substantial undertaking but is essential to the downstream success of any benchmark dataset. Unfortunately, this type of work is often undervalued and under-incentivized (Gero et al., 2023). To incentivize and recognize these efforts, datasets could be treated as citable intellectual contributions through persistent identifiers, such as a Digital Object Identifier (DOI) (Longjohn et al., 2024).
@@ 5.1 Governance and Administration
A third-party model involving a neutral entity like an international research institute or data trust presents a distinct advantage in this regard (Paprica et al., 2020). Such an organization is not constrained by jurisdictional borders and could therefore aggregate data across multiple states, provinces, and nations. This transnational approach would enable a far more comprehensive dataset.
@@ 5.2 Dataset Size
The use of synthetic data provides a potentially constructive path to scale dataset size while avoiding issues with player privacy and commercial data sharing sensitivity (Samsel et al., 2025). By generating an artificial dataset that replicates the statistical properties of real-world data without containing any personally identifiable information, synthetic data allows for broad access for research and model testing while the source data remains secure. The viability of this approach, however, is entirely contingent upon the quality and representativeness of the original dataset from which the synthetic version is derived, as well as the assumptions of the approach used to generate the synthetic data (Chen et al., 2021; Susser et al., 2024).
@@ 5.3 Defining the Target Variable
First, proxy indicators of harm such as account closure and, more commonly, voluntary self-exclusion (VSE) are used by both industry and academic studies (e.g., Finkenwirth et al., 2021; Hopfgartner et al., 2023; Percy et al., 2016). These are routinely collected alongside the data used to make predictions, can easily delineate the at-risk group through binary classification (VSE or no VSE), and are not subject to self-report bias or measurement error. However, their reliability as indicators of at-risk status is questionable. Players can choose to self-exclude or close their accounts for various reasons beyond the experience of gambling problems, including to pre-empt harms or to close one’s online account (Catania and Griffiths, 2021; Auer and Griffiths, 2016). One strategy to mitigate this risk is to use a subset of self-excluders who have a sufficiently long track record of play with the operator, to ensure players who self-exclude relatively quickly after opening an account are not used for model training.
@@
Second, screening instruments such as the PGSI (Ferris and Wynne, 2001) or Brief Biosocial Gambling Screen (Gebauer et al., 2010) have been used as target variables in recent academic studies (Auer and Griffiths, 2023; Hopfgartner et al., 2024; Kairouz et al., 2023; Murch et al., 2023; Perrot et al., 2022) and by Focal Research Consultants (Hancock et al., 2008; Schellinck and Schrans, 2011). Self-report measures offer a more reliable indicator of risk status, although are more difficult to obtain and link with the data used to make predictions, are subject to self-report biases (particularly if administered by the operator), and require a threshold for classification as at-risk to be selected in advance (Murch et al., 2023). Additionally, the measures reflect a respondent’s overall gambling behavior, which may span multiple platforms, operators, or even jurisdictions, limiting its interpretability as a label for data from any single operator’s platform. Furthermore, because self-reported risk status can change over time in response to variations in gambling engagement, repeated administration of screening instruments would improve datasets. Longitudinal measurement of the target variable would allow models to be trained and evaluated on their ability to predict transitions into or out of at-risk status, rather than relying solely on cross-sectional classification.
@@
Third, a final approach may be the use of single indicators such as loss chasing or substantial variations from normal play, which are themselves sometimes used to classify consumers as at-risk in the absence of ML (Delfabbro et al., 2024; McAuliffe et al., 2022). Several researchers have attempted to develop thresholds for standard behavioral indicators (e.g., frequency of gambling, amount spent) that differentiate lower- from higher-risk gambling to form lower risk gambling limits (Hodgins et al., 2023; Louderback et al., 2021) – these, or similar threshold-style indicators, could plausibly serve as targets, either individually or in combination.
@@
Overall, the variation and lack of a standard approach to classification presents a challenge to constructing and comparing ML models, as there is no agreed-upon specification to label training data appropriately. Given the diversity of outcome variables and definitions of risk, we do not propose a single ground truth. Instead, we recommend a suite of benchmarks that can evolve over time as the domain continues to develop and as a clearer definition of “at-risk” emerges. We do, however, advocate for transparency in how at-risk status is determined to ensure the process can be replicated in real-world environments and to help acknowledge the biases inherent in different approaches. For example, human-based labeling of cases based on opaque reasoning would not satisfy this criterion, but labels determined using a clear and reproducible set of steps would.
@@ 5.4 Demographic and Jurisdictional Diversity
Another challenge is significant diversity in player populations and jurisdictions. A successful benchmarking framework must avoid over-representing datasets from one demographic group (e.g., all men) or market (e.g., North America), except where the benchmark results are explicitly only applied within those settings. Over-representing certain demographic groups in benchmark datasets risks validating the performance of models that perform poorly in real-world settings with more diverse populations and could potentially favor biased models that disproportionately target or ignore minority groups (see Murch et al., 2024a).
There is evidence that ML models can effectively detect at-risk gambling in datasets from countries not included in training data, although training on local data can improve performance (Hopfgartner et al., 2024). A fairer benchmarking framework will offer datasets from a variety of regions to allow contributors to demonstrate the performance of their model(s) in preferred settings (e.g., EU vs. North America).
@@ 5.5 Dataset Characteristics
Jurisdictional variation may also impact the ease with which models are applied due to differences in feature availability. Additionally, differences in available products and other policies across jurisdictions will affect model applicability. This may skew the composition and relevance of benchmark datasets developed in those contexts. For example, a model built using online tracking data in Australia where players can cancel their withdrawals may attribute high importance to this behavior, which would not be present in datasets from jurisdictions that have prohibited this feature (e.g., Great Britain; UK Gambling Commission, 2021). This highlights the need for clear documentation that describes the characteristics of each benchmarking dataset and idiosyncrasies resulting from regulation or product innovation.
@@
A further consideration is the temporal structure of the data. Academic studies report building models using varying timeframes of tracking data, ranging from 12 months (e.g., Murch et al., 2023; Perrot et al., 2022) to just 30 days (Hopfgartner et al., 2024). As discussed in Section 4, there is a need to provide benchmarks for different periods that span months to days. Ideally, benchmark datasets will span an extended period (e.g., 12 months) to permit varied testing on the same dataset and the use of diverse timeframe aggregations (e.g., monthly or daily behavioral averages). A benchmark for acute risk detection (e.g., 1 to 7 days) may benefit from uniquely large datasets, since these can include all active and newly registered customers, not only those with extensive historical data.